On October 1, 2019, Gandalf, Inc. borrowed $330,000 to finance the purchase of
a building. The terms of the mortgage require payments of $2,000 to be made at
the end of every month with the first payment being due on October 31, 2019. The
length of the mortgage is thirty years, and the mortgage carries an annual interest
rate of 6%. The amount of interest expense that Gandalf would report on its 2019
income statement related to this mortgage is equal to: multiple choice

A. $4,683.46

B. $4,848.48

C. $4,950.00

D. $4,755.85

E. $4,944.74
Answer #1

Month Oct Nov Dec Total Opening Balance Monthly Interest Monthly Repayment Closing Balance c=b*0.5% e=b+c-d $ 3,30,000.00 $ 1